Statistician



How many liters of water must be mixed with 20 liters of a 90% alcohol mixture so that the alcohol content of the mixture is reduced to 50%?




















Statistician












Answer by mananth(16949)

-------------- percent ---------------- Amount
Alcohol 90 ---------------- 20 liters
Water 0 ---------------- x liters
Total 50 ---------------- 20 + x liters

90*20+0 x=50(20 +x)
1800+0x =1000+50x
-50x=-1800+1000
-50x = -800
/ -50
x = 16 of liters Water